### v1

# Boundary control systems on a 1D spatial domainBouchra Elghazi, Birgit Jacob, and Hans Zwart

Abstract
The aim of this paper is to investigate the well-posedness of a class of boundary control and observation systems on a one dimensional spatial domain. We derive a necessary and sufficient condition characterizing the well-posedness of these systems. Furthermore, we show that the well-posedness and full control and observation implies exact controllability and exact observability. The theoretical results are illustrated using Euler-Bernoulli beam models.
keywords:
Boundary control and observation systems, well-posedness, exact controllability, exact observability, Euler-Bernoulli beam models
{MSCcodes}
47D06, 49K40, 35F05, 93B52
1 Introduction
Many physical structures rarely remain rigid when subjected to force. In fact, the modelling and analysis of flexible structures like beams, or strings play a crucial role in many engineering applications such as in robotics, aerospace, and high precision machines. These structures are typically modelled by partial differential equations (PDEs), particularly, linear time-invariant PDEs have garnered significant attention. In many applications, these PDEs are controlled and observed only at their boundaries. They are naturally modelled within the framework of boundary control and observation systems. This concept was initiated by Fattorini in the 1960s [Fattorini:68], establishing a starting point for studying linear time-invariant systems governed by PDEs with boundary control. Later researchers such as Salamon [Salamon:87, salamon_realization:88], Weiss [weiss:89, Weis:89, Wei:94, weiss1989representation, weiss1994transfer] and Curtain [curtain1997salamon] developed Fattorini’s early work to the modern theory of well-posedness, admissibility and regular linear systems.
The concept of well-posedness is fundamental in the study of boundary control and observation, as it forms the basis for further control and stability analysis. Over the years, the study of well-posedness of linear time-invariant systems has attracted considerable attention, leading to a significant amount of literature. Seminal contributions in this area include the works by Lasiecka and Triggiani [LasieckaTriggiani2000a, LasieckaTriggiani2000b] and by Staffans [Sta:05]. For further reading, we also refer to [MorrisCheng:99, MorrisCheng:03, CurtainWeiss1989, CurtZwa:20, JacZw:12, TucsnakWeiss:14, Weiss2001]. Informally speaking, well-posedness refers to the property that for every initial condition in the state space and any input function in a specified space of functions, the system has a unique state trajectory and a unique output function. Moreover, the output must belong to a specified space of functions, and both the state and the output must depend continuously on the initial state and on the input. Equivalently, it concerns the well-definedness and boundedness of the four mappings from initial state to input, initial state to final state, input to state, and input to output (also called transfer function). For boundary control and observation systems that are impedance passive, there exists a simple way of characterizing the well-posedness. In fact, if the system is internally well-posed, then the well-posedness is equivalent to the boundedness of the transfer function on a vertical line in the open right half plane. Notable examples within this class are port-Hamiltonian systems which provide a powerful framework for modelling as it views physical systems from an energy-based perspective [Schaft:06, SchJelt:14, SchMas:02].
The primary focus of this article is to investigate the well-posedness of the following class of linear time-invariant systems on a one-dimensional spatial domain
(1) $\frac{\partial x}{\partial t}(\zeta,t)$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}(\zeta)\right)\mathcal{H}(\zeta)x(\zeta,t),\quad t>0,\quad\zeta\in[0,1]$
$u(t)$ $=W_{B,1}\tau(\mathcal{H}x)(t),\quad 0=W_{B,2}\tau(\mathcal{H}x)(t),\quad t>0$
$y(t)$ $=W_{C}\tau(\mathcal{H}x)(t),\quad t>0$
$x(\zeta,0)$ $=x_{0}(\zeta),\quad\zeta\in[0,1],$
where $x(\zeta,t)\in\mathbb{F}^{n}$ ($\mathbb{F}=\mathbb{R}$ or $\mathbb{C}$), $u(t),\,y(t)\in\mathbb{F}^{m}$ and $\tau$ is the trace operator, given by
$\tau(x)=\begin{pmatrix}x(1)&x^{\prime}(1)&x(0)&x^{\prime}(0)\end{pmatrix}^{\top}$.
The state space is $X=\mathrm{L^{2}}((0,1);\mathbb{F}^{n})$ with energy inner product
(2) $\left\langle f,g\right\rangle_{X}=\frac{1}{2}\int_{0}^{1}g(\zeta)^{\ast}\mathcal{H}(\zeta)f(\zeta)d\zeta.$
For simplicity, we restrict attention to the interval $[0,1]$, the same results hold for an arbitrary compact interval.
Using the port-Hamiltonian approach, internal well-posedness [GorZwaMas:2005, Villegas:07], stability and stabilizability [Aug:16, Aug:15, AugJac:14, schmid2021stabilization], observer design [LeGorrecToledoRamirezWu:23] and robust output regulation [HumLassi:18, humaloja2016robust, PaunLeGorrecRamirez:21] of the system (1) have been investigated. However, less is known on well-posedness. If $P_{2}=0$ and $P_{1}$ is invertible it has been shown that internal well-posedness, that is the operator
$A_{0}x$ $=P_{1}\frac{\partial}{\partial\zeta}\mathcal{H}x+P_{0}\mathcal{H}x,$
$D(A_{0})$ $=\left\{x\in X\,|\,\mathcal{H}x\in\mathrm{H}^{1}((0,1);\mathbb{F}^{n}),\,\begin{bmatrix}W_{B,1}\\ W_{B,2}\end{bmatrix}\tau(\mathcal{H}x)=0\right\},$
generates a $C_{0}$-semigroup on $X$, is equivalent to the well-posedness of the system, see [JacZw:12, ZwGorMasVill:10]. This class of systems covers the wave equation, the transport equation, and the Timoshenko beam equation. However, this result does not extend to the system (1), that models for example the Schrödinger equation or the Euler-Bernoulli beam equation, see the counterexample in [ElghJacZw:25]. In [ElghJacZw:25], the well-posedness of the system (1) was proven under the assumptions that $P_{1}=0$ and $\mathcal{H}$ is constant. In this paper, we extend the analysis to the general cas and we derive a simple verifiable conditions for well-posedness of the system (1).
We proceed as follows. In the next section, we begin by providing the necessary mathematical background on boundary control and observation systems and on well-posedness. In Section 3, the main result on well-posedness is presented, and followed by Section 4 which is devoted to the exact controllability and observability. The application for Euler-Bernoulli beam models are presented in Section 5. Finally, some concluding and further works are given in Section 6.
Notation. Let $\mathbb{F}=\mathbb{C}$, with $\|\cdot\|$ denoting the Euclidean norm of $\mathbb{F}^{n}$. The state of the system at time $t$ and spatial position $\zeta$ is represented by $x(\zeta,t)$, the derivative w.r.t. time $t$ is denoted as $\dot{x}$, while $x^{\prime}$ refers to the derivative of $x$ w.r.t. $\zeta$. Moreover, $u(t)\in U$ and $y(t)\in Y$ denote the inputs and the outputs, respectively. For simplicity, we write $\mathcal{H}(\zeta)x(\zeta,t)$ as $(\mathcal{H}x)(\zeta,t)$. Throughout, we denote by $\mathbb{C}^{+}_{\alpha}$ the right half-plane $\left\{s\in\mathbb{C}~|\,\mathrm{Re}(s)>\alpha\right\}$ and the symbol $\lesssim$ is used to indicate an inequality up to a constant multiple $c>0$, where $c$ is generic and may vary from line to line. For a linear (unbounded) operator $A$ on a Hilbert space $X$ with domain $D(A)$, we denote by $\rho(A)$ its resolvent set and by $\ker A$ its kernel. If $A$ is a densely defined linear operator, then its adjoint is denoted by $A^{\ast}:D(A^{\ast})\subseteq X\to X$, see [engel2000one, pazy2012semigroups] for a formal definition. For Lebesgue and Sobolev spaces, we adopt the standard notation, such as in [adams]. The space of bounded, linear operators between two Hilbert spaces $X$ to $Z$ is denoted by $\mathcal{L}(X,Z)$, while $C^{k}(X;Z)$ is the space of all functions on $X$ mapping into $Z$ that are $k$-times continuously differentiable. Henceforth, we abbreviate $\mathcal{L}(X):=\mathcal{L}(X,X)$. Additionally, for a positive bounded operator $A$, $A^{\frac{1}{2}}$ will represents the square root of $A$, see [rudin1991functional] for a formal definition.
2 Boundary control and observation systems
In this section, we provide some background and useful preliminary results on boundary control and boundary observation systems of the form
(3) $\dot{x}(t)$ $=\mathcal{A}x(t),\quad x(0)=x_{0},$
$u(t)$ $=\mathcal{B}x(t),$
$y(t)$ $=\mathcal{C}x(t).$
We briefly review some important definitions and facts based on previous work in the literature [CurtZwa:20, JacZw:12, Sta:05, tucsnak2009observation].
Definition 2.1.
[CurtZwa:20, Definition 10.1.2] Let $X$, $U$ and $Y$ be complex Hilbert spaces. We call (3) a boundary control and observation system on $(X,U,Y)$ if the following properties hold
  1. 1.
    $\mathcal{A}:D(\mathcal{A})\subset X\to X$, $\mathcal{B}:D(\mathcal{B})\subset X\to U$, $\mathcal{C}:D(\mathcal{A})\subset X\to Y$ are linear operators with $D(\mathcal{A})\subseteq D(\mathcal{B})$.
  2. 2.
    The operator $A_{0}:D(A_{0})\to X$ defined by A_0 x = Ax  for x D(A_0) := D(A) ∩ker(B) is the generator of a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $X$.
  3. 3.
    There exists an operator $B\in L(U,X)$ such that for all $u\in U$ we have $Bu\in D(\mathcal{A})$, $\mathcal{A}B\in L(U,X)$} and BBu = u,  u U.
  4. 4.
    The operator $\mathcal{C}$ is bounded from $D(A_{0})$ to $Y$. Here, $D(A_{0})$ is equipped with the graph norm.
Next, we define classical solutions for boundary control and observation systems.
Definition 2.2.
[CurtZwa:20, Definition 10.1.3] The pair $(x,y)$ is a classical solution of the boundary control and observation system (3) on $[0,T]$ if for $x_{0}\in D(\mathcal{A})$ and $u\in C^{2}([0,T];U)$, the pair $x\in C^{1}([0,T];X)$ and $y\in C([0,T];Y)$ satisfies (3) for all $t\in[0,T]$.
For sufficiently smooth inputs, it is possible to reformulate the boundary control and observation system into an abstract differential equation of the form
(4) $\dot{v}(t)$ $=A_{0}v(t)-B\dot{u}(t)+\mathcal{A}Bu(t),$
$v(0)$ $=v_{0},$
with the output defined by
(5) $y_{v}(t)=\mathcal{C}v(t).$
Since $A_{0}$ is the infinitesimal generator of a $C_{0}$-semigroup and $B$ and $\mathcal{A}B$ are bounded, the Cauchy problem (4) has a unique classical solution for $v_{0}\in D(A_{0})$ and $u\in C^{2}([0,T];U)$. Therefore, we have the following relation between the classical solutions of (3) and (4).
Theorem 2.1.
[CurtZwa:20, Theorem 10.1.4] Consider the boundary control and observation system (3) and the abstract differential equation (4) with the output (5). Assume that $u\in C^{2}([0,T];U)$. Then, if $v_{0}=x_{0}-Bu(0)\in D(A_{0})$, the classical solutions of (3) and (4)–(5) are related by
$v(t)$ $=x(t)-Bu(t),$
$y_{v}(t)$ $=y(t)-\mathcal{C}Bu(t).$
Furthermore, the classical solution of (3) is unique.
Next, we provide a formal definition of well-posedness.
Definition 1.
[JacZw:12, Definition 13.1.3] The boundary control and observation system (3) is called well-posed, if there exist $t,m>0$ such that every classical solution of (3) satisfies
(6) $\|x(t)\|^{2}+\int_{0}^{t}\|y(s)\|^{2}ds\leq m\left(\|x_{0}\|^{2}+\int_{0}^{t}\|u(s)\|^{2}ds\right).$
We note that if the inequality (6) holds for one $t>0$, then it holds for all $t>0$ [JacZw:12]. Moreover, the inequality of well-posedness is equivalent to the existence of mild solutions for an arbitrary initial condition $x_{0}\in X$ and an arbitrary input $u\in\mathrm{L}^{2}((0,t);U)$, such that $x$ is continuous and $y\in\mathrm{L}^{2}((0,t);Y)$ [JacZw:12, Section 13.1]. Further, the state and output depend continuously on the initial condition and input function.
Remark 2.2.
[JacZw:12, Section 11.1 and Theorem 11.2.1] Assume that the boundary control and observation system (3) is well-posed. If $x_{0}\in X$ and $u\in\mathrm{H}^{1}([0,T];U)$, then the mild solution of the system (3) is given by
$x(t)$ $=T(t)(x_{0}-Bu(0))+\int_{0}^{t}T(t-s)\left(\mathcal{A}Bu(s)-B\dot{u}(s)\right)ds+Bu(t).$
If, in addition, $u\in C^{2}([0,T];U)$ and $x_{0}-Bu(0)\in D(A_{0})$. Then, the classical solution is a mild solution and the corresponding output is given by
$y(t)$ $=\mathcal{C}T(t)(x_{0}-Bu(0))+\mathcal{C}\int_{0}^{t}T(t-s)\left(\mathcal{A}Bu(s)-B\dot{u}(s)\right)ds+\mathcal{C}Bu(t).$
Definition 2.
[Aug:16, Definition 3.2.12] The boundary control and observation system (3) on $(X,U,U)$ is called impedance passive if,
(7) $\mathrm{Re}\left\langle\mathcal{A}x,x\right\rangle_{X}\leq\mathrm{Re}\left\langle\mathcal{B}x,\mathcal{C}x\right\rangle_{U},\quad x\in D(\mathcal{A}).$
It is called impedance energy-preserving if
(8) $\mathrm{Re}\left\langle\mathcal{A}x,x\right\rangle_{X}=\mathrm{Re}\left\langle\mathcal{B}x,\mathcal{C}x\right\rangle_{U},\quad x\in D(\mathcal{A}).$
Definition 3.
[JacZw:12, Definition 12.1.1] Let $s\in\mathbb{C}$ and $u_{0}\in U$. The tuple $\left(u_{0}e^{st},x(t),y(t)\right)_{t\geq 0}$ is called an exponential solution of the boundary control and observation system (3) if there exist $x_{0}\in X$, $y_{0}\in Y$, such that
$\left(u_{0}e^{st},x(t),y(t)\right)=\left(u_{0}e^{st},x_{0}e^{st},y_{0}e^{st}\right),\quad\text{for a.e. }t\geq 0,$
and $(x,y)$ is the mild solution of the boundary control and observation system (3) with the input function $u(t)=u_{0}e^{st}\in C^{1}([0,t];U)$ and initial condition $x_{0}$.
Let $s\in\mathbb{C}$. If for every $u_{0}\in U$ there exists an exponential solution, and the corresponding output trajectory $y_{0}e^{st},\,t\in[0,\infty)$ is unique, then we call the mapping $u_{0}\to y_{0}$ the transfer function at $s$, denoted by $\mathcal{G}(s)$. The mapping $s\in\rho(A_{0})\to\mathcal{G}(s)$ exists for every $s\in\rho(A_{0})$ [JacZw:12].
Theorem 4.
[JacZw:12, Theorem 12.1.3] The transfer function of the boundary control and observation system (3) is given by
$\mathcal{G}(s)=\mathcal{C}(sI-A_{0})^{-1}(\mathcal{A}B-sB)+\mathcal{C}B,\quad s\in\rho(A_{0}).$
Moreover, every exponential solution of (3) is also a classical solution. Furthermore, for $s\in\rho(A_{0})$ and $u_{0}\in U$, $\mathcal{G}(s)u_{0}$ can also be calculated as the (unique) solution of
(9) $sx_{0}$ $=\mathcal{A}x_{0},$
$u_{0}$ $=\mathcal{B}x_{0},$
$\mathcal{G}(s)u_{0}$ $=\mathcal{C}x_{0}.$
with $x_{0}\in D(\mathcal{A})$. Further, $x_{0}\in D(\mathcal{A})$ is uniquely determined by (9).
Corollary 2.3.
Let $u_{0}\in U$ and $x_{0}\in D(\mathcal{A})$ be the unique solution of (9). If the boundary control and observation system is impedance passive, then we have the following inequality
(10) $\mathrm{Re}(s)\|x_{0}\|^{2}_{X}\leq\mathrm{Re}\left\langle u_{0},\mathcal{G}(s)u_{0}\right\rangle,$
for all $s$ such that $\mathcal{G}(s)$ exists.
Proof 2.4.
By Definition 3, the transfer function is related to the exponential solution via
$\left(u_{0}e^{st},x_{0}e^{st},\mathcal{G}(s)u_{0}e^{st}\right)_{t\geq 0}.$
By Theorem 4 $x_{0}\in D(\mathcal{A})$ and (9) is satisfied. The impedance passivity of the system gives
$\mathrm{Re}\left\langle\mathcal{A}x_{0}e^{st},x_{0}e^{st}\right\rangle_{X}$ $\leq\mathrm{Re}\left\langle\mathcal{B}x_{0},\mathcal{C}x_{0}\right\rangle$
$=\left\langle u_{0}e^{st},\mathcal{G}(s)u_{0}e^{st}\right\rangle.$
Using (9), we get
$\mathrm{Re}\left\langle sx_{0}e^{st},x_{0}e^{st}\right\rangle_{X}\leq\mathrm{Re}\left\langle u_{0}e^{st},\mathcal{G}(s)u_{0}e^{st}\right\rangle.$
We divide by $e^{2\mathrm{Re}(s)t}$ and obtain
$\mathrm{Re}(s)\|x_{0}\|_{X}^{2}\leq\mathrm{Re}\left\langle u_{0},\mathcal{G}(s)u_{0}\right\rangle.$
Whereas well-posedness requires the inequality (6) to be satisfied, for impedance passive systems, well-posedness is already determined by the boundedness of the transfer function on some vertical line in the open right half-plane. We have the following result.
Theorem 5.
[Sta:02, Theorem 5.1] An impedance passive boundary control and observation system is well-posed if and only if its transfer function $\mathcal{G}$ is bounded on some vertical line in the open right half plane $\mathbb{C}^{+}_{0}$, i.e. $\mathcal{G}:\mathbb{C}^{+}_{0}\longrightarrow\mathcal{L}(U)$ satisfies
$\text{There exists }r>0\text{ such that }\sup_{\omega\in\mathbb{R}}\|\mathcal{G}(r+i\omega)\|<\infty.$
For well-posed boundary control and observation systems, even if the transfer function is bounded on some right half plane, this does not imply that the limit as $\mathrm{Re}(s)\to\infty$ exists. Well-posed systems with this property are called regular.
Definition 6.
[Wei:94, Definition 4.4] Let $\mathcal{G}$ be the transfer function of a well-posed boundary control and observation system. The boundary control and observation system is called regular if $\lim_{s\in\mathbb{R},s\to\infty}\mathcal{G}(s)$ exists. If the boundary control system is regular, then the feedthrough term $D$ is defined as $D=\underset{s\in\mathbb{R},s\to\infty}{\lim}\mathcal{G}(s)$.
Next, we state that well-posed boundary control and observation systems are stable under bounded perturbations and admissible feedback operators.
Lemma 2.5.
([engel2000one, Theorem 1.3] and [tucsnak2009observation, Remark 2.11.3]) Let $\mathcal{G}$ be the transfer function of the boundary control and observation system generated by $\left(\mathcal{A},\mathcal{B},\mathcal{C}\right)$, and let $P$ be a bounded linear operator on $X$. Then the boundary control and observation system generated by $\left(\mathcal{A},\mathcal{B},\mathcal{C}\right)$ is well-posed if and only if the the boundary control and observation system generated by $\left(\mathcal{A}+P,\mathcal{B},\mathcal{C}\right)$ is well-posed. Moreover, if $\mathcal{G}_{P}$ is the transfer function of the boundary control and observation system generated by $\left(\mathcal{A}+P,\mathcal{B},\mathcal{C}\right)$, then
$\lim_{s\in\mathbb{R},s\to\infty}\mathcal{G}(s)=\lim_{ss\in\mathbb{R},s\to\infty}\mathcal{G}_{P}(s)$
and
$\lim_{\mathrm{Re}(s)\to\infty}\mathcal{G}(s)=\lim_{\mathrm{Re}(s)\to\infty}\mathcal{G}_{P}(s).$
Theorem 7.
([Wei:94, Proposition 4.9] and [JacZw:12, Theorem 13.1.12]) Let $\mathcal{G}$ be the transfer function of the well-posed boundary control and observation system
$\dot{x}(t)$ $=\mathcal{A}x(t),$
$u(t)$ $=\mathcal{B}x(t),$
$y(t)$ $=\mathcal{C}x(t).$
Let $F$ be a bounded linear operator from $Y$ to $U$ such that the inverse of $I+\mathcal{G}(s)F$ $\left(\text{or }I+F\mathcal{G}(s)\right)$ exists and is bounded in $s$ in some right half-plane. Then the closed-loop system is again well-posed, that is, the boundary control and observation system
$\dot{x}(t)$ $=\mathcal{A}x(t),$
$u(t)$ $=(\mathcal{B}+F\mathcal{C})x(t),$
$y(t)$ $=\mathcal{C}x(t),$
is again well-posed. The operator $F$ is called an admissible feedback operator.
We conclude the section by a useful lemma.
Lemma 2.6.
Let $\left(\mathcal{A},\mathcal{B}_{0},\mathcal{C}_{0}\right)$ and $\left(\mathcal{A},\mathcal{B}_{1},\mathcal{C}_{1}\right)$ be two well-posed boundary control and observation systems on $\left(X,U,Y\right)$ and let $Q\coloneqq\begin{pmatrix}Q_{11}&Q_{12}\\ Q_{21}&Q_{22}\end{pmatrix}\in\mathcal{L}(U\times Y)$ be such that
(11) $Q\begin{pmatrix}\mathcal{B}_{0}x\\ \mathcal{C}_{0}x\end{pmatrix}=\begin{pmatrix}\mathcal{B}_{1}x\\ \mathcal{C}_{1}x\end{pmatrix},\quad x\in D(\mathcal{A}).$
We denote by $\mathcal{G}_{0}$ and $\mathcal{G}_{1}$ the transfer functions of $\left(\mathcal{A},\mathcal{B}_{0},\mathcal{C}_{0}\right)$ and $\left(\mathcal{A},\mathcal{B}_{1},\mathcal{C}_{1}\right)$, respectively and we assume that $\lim_{\mathrm{Re}(s)\to\infty}\mathcal{G}_{0}(s)=0$. If $Q$ is invertible, then $Q_{11}$ is invertible.
Proof 2.7.
Since the two systems $\left(\mathcal{A},\mathcal{B}_{0},\mathcal{C}_{0}\right)$ and $\left(\mathcal{A},\mathcal{B}_{1},\mathcal{C}_{1}\right)$ are well-posed, there exist $\alpha>0$ and $M_{0},M_{1}>0$ such that
$\sup_{\mathrm{Re}(s)>\alpha}\|\mathcal{G}_{0}(s)\|$ $<M_{0},$
$\sup_{\mathrm{Re}(s)>\alpha}\|\mathcal{G}_{1}(s)\|$ $<M_{1}.$
By Theorem 4, for $u_{0}\in U$ and $s\in\mathbb{C}^{+}_{\alpha}$, $\mathcal{G}_{0}(s)u_{0}$ is the unique solution of
$sx_{0}$ $=\mathcal{A}x_{0},\quad x_{0}\in D(\mathcal{A}),$
$u_{0}$ $=\mathcal{B}_{0}x_{0},$
$\mathcal{G}_{0}(s)u_{0}$ $=\mathcal{C}_{0}x_{0}.$
Therefore, applying the matrix $\begin{pmatrix}Q_{11}&Q_{12}\\ Q_{21}&Q_{22}\end{pmatrix}$ to the input and the output equations we obtain that
$sx_{0}$ $=\mathcal{A}x_{0},$
$\begin{pmatrix}\mathcal{B}_{1}x_{0}\\ \mathcal{C}_{1}x_{0}\end{pmatrix}$ $=\begin{pmatrix}Q_{11}&Q_{12}\\ Q_{21}&Q_{22}\end{pmatrix}\begin{pmatrix}u_{0}\\ \mathcal{G}_{0}(s)u_{0}\end{pmatrix}$.
Defining
$u_{1}(s)$ $=Q_{11}u_{0}+Q_{12}\mathcal{G}_{0}(s)u_{0},$
$y_{1}(s)$ $=Q_{21}u_{0}+Q_{22}\mathcal{G}_{0}(s)u_{0},$
it follows that
$sx_{0}$ $=\mathcal{A}x_{0},$
$\mathcal{B}_{1}x_{0}$ $=u_{1}(s),$
$\mathcal{C}_{1}x_{0}$ $=y_{1}(s).$
As the transfer function is unique, we have $y_{1}(s)=\mathcal{G}_{1}(s)u_{1}(s)$. Hence, we have shown that for every $u_{0}\in U$ and $s\in\mathbb{C}^{+}_{\alpha}$ there exists $u_{1}(s)\in U$ such that
(12) $\begin{pmatrix}u_{1}(s)\\ \mathcal{G}_{1}(s)u_{1}(s)\end{pmatrix}=Q\begin{pmatrix}u_{0}\\ \mathcal{G}_{0}(s)u_{0}\end{pmatrix}$.
In addition, as $Q$ is invertible, the same argument shows that for every $u_{1}\in U$ and $s\in\mathbb{C}^{+}_{\alpha}$ there exists $u_{0}(s)\in U$ such that
(13) $Q\begin{pmatrix}u_{0}(s)\\ \mathcal{G}_{0}(s)u_{0}(s)\end{pmatrix}=\begin{pmatrix}u_{1}\\ \mathcal{G}_{1}(s)u_{1}\end{pmatrix}$.
Using the equation (12), for $s\in\mathbb{C}^{+}_{\alpha}$ and $u_{0}\in U$ we have
$M_{1}\|Q_{11}u_{0}\|+M_{1}\|Q_{12}\mathcal{G}_{0}(s)u_{0}\|$ $\geq M_{1}\|Q_{11}u_{0}+Q_{12}\mathcal{G}_{0}(s)u_{0}\|$
$\geq\|\mathcal{G}_{1}(s)u_{1}(s)\|$
$=\|Q_{21}u_{0}+Q_{22}\mathcal{G}_{0}(s)u_{0}\|$
$\geq\|Q_{21}u_{0}\|-\|Q_{22}\mathcal{G}_{0}(s)u_{0}\|.$
As the system $\left(\mathcal{A},\mathcal{B}_{0},\mathcal{C}_{0}\right)$ is regular with a feedthrough 0, taking the limit as $s\to\infty$}, we get that
(14) $M_{1}\|Q_{11}u_{0}\|\geq\|Q_{21}u_{0}\|,\quad u_{0}\in U.$
Since $Q$ is bounded and invertible, the inverse $Q^{-1}$ is bounded as well. Thus $Q$ is bounded from below, i.e. there exists $C>0$ such that for $\begin{pmatrix}u_{0}\\ y_{0}\end{pmatrix}\in U\times Y$ we have
$\left\|Q\begin{pmatrix}u_{0}\\ y_{0}\end{pmatrix}\right\|\geq C\left\|\begin{pmatrix}u_{0}\\ y_{0}\end{pmatrix}\right\|$.
However, by (14) for every $u_{0}\in U$
$\left(1+M_{1}^{2}\right)\|Q_{11}u_{0}\|^{2}$ $\geq\|Q_{11}u_{0}\|^{2}+\|Q_{21}u_{0}\|^{2}$
$=\left\|Q\begin{pmatrix}u_{0}\\ 0\end{pmatrix}\right\|^{2}$
$\geq C^{2}\left\|\begin{pmatrix}u_{0}\\ 0\end{pmatrix}\right\|^{2}=C^{2}\|u_{0}\|^{2}.$
Thus, for $\tilde{C}=\frac{C^{2}}{1+M_{1}^{2}}$,
$\|Q_{11}u_{0}\|^{2}\geq\tilde{C}\|u_{0}\|^{2}.$
This implies that $\ker Q_{11}=\{0\}$ and $\operatorname{ran}Q_{11}$ is closed.
Assuming $Q_{11}$ is not surjective, there exists $u_{1}\in U$ with $u_{1}\not=0$ and $u_{1}\in\operatorname{ran}Q_{11}^{\perp}$. Using equation (13), we have that for $s\in\mathbb{C}^{+}_{\alpha}$ there exists $u_{0}(s)\in U$ such that
(15) $\begin{pmatrix}Q_{11}&Q_{12}\\ Q_{21}&Q_{22}\end{pmatrix}\begin{pmatrix}u_{0}(s)\\ \mathcal{G}_{0}(s)u_{0}(s)\end{pmatrix}=\begin{pmatrix}u_{1}\\ \mathcal{G}_{1}(s)u_{1}\end{pmatrix}$.
Since $Q$ is invertible and $\mathcal{G}_{1}(s)$ is bounded on $\mathbb{C}^{+}_{\alpha}$, (15) gives that $u_{0}(s)$ is bounded on $\mathbb{C}^{+}_{\alpha}$ as well. Now,
$\|u_{1}\|^{2}$ $=\displaystyle\lim_{s\in\mathbb{R},s\to\infty}\|u_{1}\|^{2}$
$=\displaystyle\lim_{s\in\mathbb{R},s\to\infty}\left\langle\begin{pmatrix}u_{1}\\ 0\end{pmatrix},Q\begin{pmatrix}u_{0}(s)\\ \mathcal{G}_{0}(s)u_{0}(s)\end{pmatrix}\right\rangle$
$=\displaystyle\lim_{s\in\mathbb{R},s\to\infty}\left\langle u_{1},\left(Q_{11}+Q_{12}\mathcal{G}_{0}(s)\right)u_{0}(s)\right\rangle$
$=\displaystyle\lim_{s\in\mathbb{R},s\to\infty}\left\langle u_{1},Q_{12}\mathcal{G}_{0}(s)u_{0}(s)\right\rangle$
$=0$
which gives a contradiction. We conclude that $Q_{11}$ is invertible.
3 Analysis of well-posedness
We now turn our attention to the class of systems
(16) $\frac{\partial x}{\partial t}(\zeta,t)$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}(\zeta)\right)\mathcal{H}(\zeta)x(\zeta,t),\quad t>0,\quad\zeta\in[0,1]$
$u(t)$ $=W_{B,1}\tau(\mathcal{H}x)(t),\quad 0=W_{B,2}\tau(\mathcal{H}x)$
$y(t)$ $=W_{C}\tau(\mathcal{H}x)(t)$
$x(\zeta,0)$ $=x_{0}(\zeta),\quad\zeta\in[0,1]$
where $\tau:\mathrm{H}^{2}((0,1);\mathbb{F}^{n})\to\mathbb{F}^{4n}$ is the trace operator given by
$\tau(x)=\begin{pmatrix}x(1)&x^{\prime}(1)&x(0)&x^{\prime}(1)\end{pmatrix}^{\top}$.
We make the following assumptions.
Assumption 3.1.
  • $P_{1},\,P_{2}\in\mathbb{F}^{n\times n}$ such that $P_{1}$ is self-adjoint and $P_{2}$ is skew-adjoint and invertible;
  • $P_{0}\in\mathrm{L}^{\infty}((0,1);\mathbb{F}^{n\times n})$;
  • $\mathcal{H}\in C^{1}([0,1];\mathbb{F}^{n\times n})$ such that $\mathcal{H}(\zeta)$ is self-adjoint for all $\zeta\in[0,1]$ and there exist $M,m>0$ such that $mI\leq\mathcal{H}(\zeta)\leq MI$ for all $\zeta\in[0,1]$;
  • The matrices $W_{B,1},W_{C}\in\mathbb{F}^{m\times 4n}$ with $0<m\leq 2n$, and $W_{B,2}\in\mathbb{F}^{(2n-m)\times 4n}$ satisfying that $\left[\begin{smallmatrix}W_{B,1}\\ W_{B,2}\\ W_{C}\end{smallmatrix}\right]$ has full row rank.
We choose $X=\mathrm{L}^{2}((0,1);\mathbb{F}^{n})$ equipped with the inner product
(17) $\left\langle f,g\right\rangle_{X}=\frac{1}{2}\int_{0}^{1}g(\zeta)^{\ast}\mathcal{H}(\zeta)f(\zeta)d\zeta.$
We note that the standard $\mathrm{L^{2}}$-norm is equivalent to the norm defined by (17), i.e., for all $f\in\mathrm{L^{2}}((0,1);\mathbb{F}^{n})$
$\frac{m}{2}\|f\|^{2}_{\mathrm{L^{2}}}\leq\|f\|^{2}_{X}\leq\frac{M}{2}\|f\|^{2}_{\mathrm{L^{2}}}.$
Remark 3.1.
By defining
(18) $\mathcal{A}x$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}\right)\mathcal{H}x,$
$\mathcal{B}x$ $=W_{B,1}\tau(\mathcal{H}x),$
$\mathcal{C}x$ $=W_{C}\tau(\mathcal{H}x),$
$D(\mathcal{A})$ $=\left\{x\in\mathrm{L}^{2}((0,1);\mathbb{F}^{n})\,|\,\mathcal{H}x\in\mathrm{H}^{2}((0,1);\mathbb{F}^{n}),\,W_{B,2}\tau(\mathcal{H}x)=0\right\},$
$D(\mathcal{B})$ $=D(\mathcal{A}),$
the system (16) is a boundary control and observation system, provided the operator $A_{0}\coloneqq\mathcal{A}_{\mid\ker\mathcal{B}}$ is the infinitesimal generator of a $C_{0}$-semigroup on $X$, see [GorZwaMas:2005].
Remark 3.2.
If (16) is an impedance passive boundary control and observation system, then the operator $A_{0}=\mathcal{A}_{\mid\ker\mathcal{B}}$ is dissipative and in this case the $C_{0}$-semigroup generated by $A_{0}$ is a contraction semigroup on $X$, see [Aug:16, Villegas:07]. If in addition, $m=2n$ and $P_{0}^{\ast}(\zeta)=-P_{0}(\zeta)$ for a.e. $\zeta\in(0,1)$, then the impedance passivity is equivalent to
$\begin{bmatrix}\tilde{W}_{B}\Sigma\tilde{W}_{B}^{\ast}&\tilde{W}_{B}\Sigma\tilde{W}_{C}^{\ast}\\ \tilde{W}_{C}\Sigma\tilde{W}_{B}^{\ast}&\tilde{W}_{C}\Sigma\tilde{W}_{C}^{\ast}\end{bmatrix}^{-1}\leq\begin{bmatrix}0&I\\ I&0\end{bmatrix}$,
and impedance energy preserving is equivalent to
$\begin{bmatrix}\tilde{W}_{B}\Sigma\tilde{W}_{B}^{\ast}&\tilde{W}_{B}\Sigma\tilde{W}_{C}^{\ast}\\ \tilde{W}_{C}\Sigma\tilde{W}_{B}^{\ast}&\tilde{W}_{C}\Sigma\tilde{W}_{C}^{\ast}\end{bmatrix}^{-1}=\begin{bmatrix}0&I\\ I&0\end{bmatrix}$,
where
$\tilde{W}_{B}$ $=\frac{1}{\sqrt{2}}\left[\begin{matrix}W_{B,1}\\ W_{B,2}\end{matrix}\right]\left[\begin{matrix}R&I\\ -R&I\end{matrix}\right],\quad\tilde{W}_{C}=\frac{1}{\sqrt{2}}W_{C}\left[\begin{matrix}R&I\\ -R&I\end{matrix}\right],$
$\Sigma$ $=\left[\begin{matrix}0&I\\ I&0\end{matrix}\right],\quad R=\left[\begin{matrix}0&-P_{2}^{-1}\\ P_{2}^{-1}&P_{2}^{-1}P_{1}P_{2}^{-1}\end{matrix}\right],$
see [AugJac:14, Aug:16]. Furthermore, by Theorem 2.1, for smooth input and initial conditions, the system (16) has classical solutions over the time interval $[0,T]$, $T>0$. Additionally, the output $y$ is well-defined and continuous on this interval. Well-posedness, on the other hand, not only guarantees the existence of mild solutions but also ensures that the output function $y$ satisfies $y\in\mathrm{L}^{2}((0,T);\mathbb{F}^{m})$ for arbitrary initial conditions $x_{0}\in X$ and inputs $u\in\mathrm{L}^{2}((0,T);\mathbb{F}^{m})$.
In what follows, we define
(19) $\mathcal{A}_{e}x$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}\right)\mathcal{H}x,$
$\mathcal{B}_{e}x$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}P_{2}\frac{\partial}{\partial\zeta}(\mathcal{H}x)(1)+\tfrac{1}{2}P_{1}(\mathcal{H}x)(1)\\ P_{2}\frac{\partial}{\partial\zeta}(\mathcal{H}x)(0)+\frac{1}{2}P_{1}(\mathcal{H}x)(0)\end{bmatrix}$,
$\mathcal{C}_{e}x$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}(\mathcal{H}x)(1)\\ -(\mathcal{H}x)(0)\end{bmatrix}$,
$D(\mathcal{A}_{e})$ $=\left\{x\in\mathrm{L}^{2}((0,1);\mathbb{F}^{n})\,|\,\mathcal{H}x\in\mathrm{H}^{2}((0,1);\mathbb{F}^{n})\right\},$
$D(\mathcal{B}_{e})$ $=D(\mathcal{A}_{e}).$
We now proceed with the analysis of the well-posedness of the system (16). As an initial step, we state the following proposition.
Proposition 3.3.
With the operators in (19), consider the boundary control and observation system
(20) $\dot{x}(t)$ $=\mathcal{A}_{e}x(t),\quad x(0)=x_{0},$
$u_{e}(t)$ $=\mathcal{B}_{e}x(t),$
$y_{e}(t)$ $=\mathcal{C}_{e}x(t),\quad t>0$
with the state $x$, the input $u_{e}$ and the output $y_{e}$. We assume that the system satisfies the conditions of Assumption 3.1. Then, the system (20) is well-posed and regular. In particular, for some $\alpha>0$ and $c>0$, the transfer function $\mathcal{G}_{e}(s)$ satisfies
$\|\mathcal{G}_{e}(s)\|_{\mathcal{L}(U)}\leq\frac{c}{\mathrm{Re}(s)},\quad s\in\mathbb{C}^{+}_{\alpha}$.
Moreover, if $P_{0}(\zeta)^{\ast}=-P_{0}(\zeta)$ for a.e. $\zeta\in(0,1)$, then the system (20) is impedance energy preserving.
Proof 3.4.
Using integration by parts, $P_{2}^{\ast}=-P_{2}$, $P_{1}^{\ast}=P_{1}$ and $\mathcal{H}(\zeta)^{\ast}=\mathcal{H}(\zeta)$ for all $\zeta\in[0,1]$, we obtain that for $x\in D(\mathcal{A}_{e})$
$\mathrm{Re}\left\langle\mathcal{A}_{e}x,x\right\rangle_{X}$ $=\mathrm{Re}\left\langle P_{2}\frac{d^{2}}{d\zeta^{2}}\mathcal{H}x+P_{1}\frac{d}{d\zeta}\mathcal{H}x+P_{0}\mathcal{H}x,x\right\rangle_{X}$
$=\frac{1}{2}\mathrm{Re}\left\langle P_{2}\frac{d^{2}}{d\zeta^{2}}\mathcal{H}x,\mathcal{H}x\right\rangle_{\mathrm{L}^{2}}+\mathrm{Re}\left\langle P_{1}\frac{d}{d\zeta}\mathcal{H}x,\mathcal{H}x\right\rangle_{\mathrm{L}^{2}}+\left\langle P_{0}\mathcal{H}x,\mathcal{H}x\right\rangle_{\mathrm{L}^{2}}$
$=\frac{1}{2}\mathrm{Re}\left[(\mathcal{H}x)(\zeta)^{\ast}\left(P_{2}\frac{d}{d\zeta}(\mathcal{H}x)(\zeta)+\frac{1}{2}P_{1}(\mathcal{H}x)(\zeta)\right)\right]_{0}^{1}+\left\langle P_{0}\mathcal{H}x,\mathcal{H}x\right\rangle_{\mathrm{L}^{2}}$
$=\mathrm{Re}\left\langle u_{e},y_{e}\right\rangle_{\mathbb{F}^{2n}}+\left\langle P_{0}\mathcal{H}x,\mathcal{H}x\right\rangle_{\mathrm{L}^{2}}.$
We see that if $P_{0}(\zeta)^{\ast}=-P_{0}(\zeta)$ for a.e. $\zeta\in(0,1)$, then $\left\langle P_{0}\mathcal{H}x,\mathcal{H}x\right\rangle_{\mathrm{L}^{2}}=0$. Thus, the system is impedance energy preserving. Nonetheless, as $P_{0}\mathcal{H}$, seen as a multiplication operator on $X$, is a bounded operator on the state space $X$, the well-posedness and regularity remains unchanged under bounded perturbations, see Lemma 2.5. Moreover, the feedthrough term will remain the same. Thus, without loss of generality, we assume that $P_{0}=0$. Therefore, the system is impedance passive. By Remark 3.2, $A_{e}\coloneqq\mathcal{A}_{e\mid\ker\mathcal{B}_{0}}$ generates a contraction semigroup on $X$. Next, we show that the system is well-posed. Since $A_{e}$ generates a contraction semigroup, by Theorem 5 it remains to prove that the transfer function of the system is bounded on some vertical line in the open right half-plane. As both $u_{e}$ and $y_{e}$ in (19) contain the common factor $\frac{1}{2}$, we may, without loss of generality, remove this factor, as boundedness is preserved under multiplication by a nonzero constant.
Let $s=r+i\omega\in\mathbb{C}$, $r>1$ and $u_{e,0}\in\mathbb{F}^{m}$. By $x_{0}$ we denote the solution of the ordinary differential equation
(21) $sx_{0}(\zeta)=P_{2}(\mathcal{H}x_{0})^{\prime\prime}(\zeta)+P_{1}(\mathcal{H}x_{0})^{\prime}(\zeta),$
with the input $u_{e,0}$
(22) $u_{e,0}=\begin{bmatrix}P_{2}(\mathcal{H}x_{0})^{\prime}(1)+\frac{1}{2}P_{1}(\mathcal{H}x_{0})(1)\\ P_{2}(\mathcal{H}x_{0})^{\prime}(0)+\frac{1}{2}P_{1}(\mathcal{H}x_{0})(0)\end{bmatrix}=\begin{bmatrix}u_{e,0}^{1}\\ u_{e,0}^{2}\end{bmatrix}$,
and we define the output $y_{e,0}$ by
(23) $y_{e,0}\coloneqq\begin{bmatrix}(\mathcal{H}x_{0})(1)\\ -(\mathcal{H}x_{0})(0)\end{bmatrix}=\begin{bmatrix}y_{e,0}^{1}\\ y_{e,0}^{2}\end{bmatrix}$.
We note that for every $\zeta\in[0,1]$, $mI\leq H(\zeta)\leq MI$, the norms $\|\mathcal{H}(\zeta)\cdot\|$, $\|\mathcal{H}^{\frac{1}{2}}(\zeta)\cdot\|$ and the Euclidean norm on $\mathbb{F}^{n}$ are equivalent and the equivalence constant can be chosen independent of $\zeta$. Using this and the Fundamental Theorem of Calculus, we have
$\|y_{e,0}\|^{2}$ $=\|(\mathcal{H}x_{0})(1)\|^{2}+\|(\mathcal{H}x_{0})(0)\|^{2}$
$\lesssim x_{0}(1)^{\ast}\mathcal{H}(1)x_{0}(1)+x_{0}(0)^{\ast}\mathcal{H}(0)x_{0}(0)$
$=\ \int_{0}^{1}\frac{d}{d\zeta}\left((2\zeta-1)x_{0}(\zeta)^{\ast}\mathcal{H}(\zeta)x_{0}(\zeta)\right)d\zeta$
= $=\ 2\int_{0}^{1}x_{0}(\zeta)^{\ast}\mathcal{H}(\zeta)x_{0}(\zeta)d\zeta+\int_{0}^{1}(2\zeta-1)\left((\mathcal{H}x_{0})^{\prime}(\zeta)\right)^{\ast}x_{0}(\zeta)d\zeta$
$\qquad\quad+\int_{0}^{1}(2\zeta-1)(\mathcal{H}x_{0})(\zeta)^{\ast}x_{0}(\zeta)^{\prime}d\zeta$
= $=\ 4\|x_{0}\|_{X}^{2}+\int_{0}^{1}(2\zeta-1)\left[\left((\mathcal{H}x_{0})^{\prime}(\zeta)\right)^{\ast}x_{0}(\zeta)+x_{0}(\zeta)^{\ast}(\mathcal{H}x_{0})^{\prime}(\zeta)\right]d\zeta$
$\qquad\quad-\int_{0}^{1}(2\zeta-1)x_{0}(\zeta)^{\ast}\mathcal{H}(\zeta)^{\prime}x_{0}(\zeta)d\zeta$
= $=\ 4\|x_{0}\|_{X}^{2}-\int_{0}^{1}(2\zeta-1)x_{0}(\zeta)^{\ast}\mathcal{H}(\zeta)^{\prime}x_{0}(\zeta)d\zeta$
$\qquad\quad+2\int_{0}^{1}(2\zeta-1)\mathrm{Re}\left[x_{0}(\zeta)^{\ast}(\mathcal{H}x_{0})^{\prime}(\zeta)\right]d\zeta.$
Since $\mathcal{H}\in C^{1}([0,1];\mathbb{F}^{n})$, we obtain
(24) $\|y_{e,0}\|^{2}\lesssim\|x_{0}\|_{X}^{2}+2\int_{0}^{1}(2\zeta-1)\mathrm{Re}\left[x_{0}(\zeta)^{\ast}(\mathcal{H}x_{0})^{\prime}(\zeta)\right]d\zeta.$
For simplicity, we call $h_{0}=\mathcal{H}x_{0}$. Using equation (21), we calculate
$h_{0}^{\prime}(\zeta)^{\ast}$ $x_{0}(\zeta)+x_{0}(\zeta)^{\ast}h_{0}^{\prime}(\zeta)$
$=\frac{1}{s}h_{0}^{\prime}(\zeta)^{\ast}\left(P_{2}h_{0}^{\prime\prime}(\zeta)+P_{1}h_{0}^{\prime}(\zeta)\right)+\frac{1}{\overline{s}}\left(P_{2}h_{0}^{\prime\prime}(\zeta)+P_{1}h_{0}^{\prime}(\zeta)\right)^{\ast}h_{0}^{\prime}(\zeta)$
$=\frac{\overline{s}}{|s|^{2}}h_{0}^{\prime}(\zeta)^{\ast}\left(P_{2}h_{0}^{\prime\prime}(\zeta)+P_{1}h_{0}^{\prime}(\zeta)\right)+\frac{s}{|s|^{2}}\left(P_{2}h_{0}^{\prime\prime}(\zeta)+P_{1}h_{0}^{\prime}(\zeta)\right)^{\ast}h_{0}^{\prime}(\zeta)$
$=\frac{r}{|s|^{2}}\left[h_{0}^{\prime}(\zeta)^{\ast}\left(P_{2}h_{0}^{\prime\prime}(\zeta)+P_{1}h_{0}^{\prime}(\zeta)\right)+\left(P_{2}h_{0}^{\prime\prime}(\zeta)+P_{1}h_{0}^{\prime}(\zeta)\right)^{\ast}h_{0}^{\prime}(\zeta)\right]$
$\qquad-\frac{i\omega}{|s|^{2}}\left[h_{0}^{\prime}(\zeta)^{\ast}\left(P_{2}h_{0}^{\prime\prime}(\zeta)+P_{1}h_{0}^{\prime}(\zeta)\right)-\left(P_{2}h_{0}^{\prime\prime}(\zeta)+P_{1}h_{0}^{\prime}(\zeta)\right)^{\ast}h_{0}^{\prime}(\zeta)\right]$
$=\frac{r}{|s|^{2}}\left[h_{0}^{\prime}(\zeta)^{\ast}sx_{0}(\zeta)+\overline{s}x_{0}(\zeta)^{\ast}h_{0}^{\prime}(\zeta)\right]-\frac{2i\omega}{|s|^{2}}\mathrm{Re}\left[h_{0}^{\prime}(\zeta)^{\ast}P_{2}h_{0}^{\prime\prime}(\zeta)\right]$
$=\frac{2r}{|s|^{2}}\mathrm{Re}\left[sh_{0}^{\prime}(\zeta)^{\ast}x_{0}(\zeta)\right]-\frac{i\omega}{|s|^{2}}\left(\left(h_{0}^{\prime}(\zeta)\right)^{\ast}P_{2}h_{0}^{\prime}(\zeta)\right)^{\prime}.$
Substituting the previous expression into (24) and using Cauchy-Schwarz inequality, the inequality (24) becomes
$\|y_{e,0}\|^{2}$ $\lesssim\|x_{0}\|_{X}^{2}+2\int_{0}^{1}(2\zeta-1)\mathrm{Re}\left[x_{0}(\zeta)^{\ast}h_{0}^{\prime}(\zeta)\right]d\zeta$
$=\|x_{0}\|_{X}^{2}+\frac{2r}{|s|^{2}}\int_{0}^{1}(2\zeta-1)\mathrm{Re}\left[sh_{0}^{\prime}(\zeta)^{\ast}x_{0}(\zeta)\right]d\zeta$
$\qquad-\frac{i\omega}{|s|^{2}}\int_{0}^{1}(2\zeta-1)\left[h_{0}^{\prime}(\zeta)^{\ast}P_{2}h_{0}^{\prime}(\zeta)\right]^{\prime}d\zeta$
$\lesssim\|x_{0}\|_{X}^{2}+\frac{r|s|}{|s|^{2}}\int_{0}^{1}\|h_{0}^{\prime}(\zeta)\|\|x_{0}(\zeta)\|d\zeta-\frac{i\omega}{|s|^{2}}\int_{0}^{1}(2\zeta-1)\left[h_{0}^{\prime}(\zeta)^{\ast}P_{2}h_{0}^{\prime}(\zeta)\right]^{\prime}d\zeta$
$\lesssim\|x_{0}\|_{X}^{2}+\frac{r}{|s|}\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}\|x_{0}\|_{\mathrm{L^{2}}}-\frac{i\omega}{|s|^{2}}\Big[(2\zeta-1)h_{0}^{\prime}(\zeta)^{\ast}P_{2}h_{0}^{\prime}(\zeta)\Big]_{0}^{1}$
$\qquad+\frac{2i\omega}{|s|^{2}}\int_{0}^{1}h_{0}^{\prime}(\zeta)^{\ast}P_{2}h_{0}^{\prime}(\zeta)d\zeta.$
Using Cauchy-Schwarz inequality and the fact that the norms $\|\cdot\|_{\mathrm{L}^{2}}$ and $\|\cdot\|_{X}$ are equivalent, we obtain that
(25) $\|y_{e,0}\|^{2}\lesssim\|x_{0}\|_{X}^{2}+\frac{r}{|s|}\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}\|x_{0}\|_{X}-\frac{i\omega}{|s|^{2}}\Big[(2\zeta-1)h_{0}^{\prime}(\zeta)^{\ast}P_{2}h_{0}^{\prime}(\zeta)\Big]_{0}^{1}+\frac{1}{|s|}\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}^{2}.$
On the other hand, using the equalities (22) and (23), together with the invertibility of the matrix $P_{2}$, we have that
$\Big[(2$ $\zeta-1)h_{0}^{\prime}(\zeta)^{\ast}P_{2}h_{0}^{\prime}(\zeta)\Big]_{0}^{1}$
$=h_{0}^{\prime}(1)^{\ast}P_{2}h_{0}^{\prime}(1)+h_{0}^{\prime}(0)^{\ast}P_{2}h_{0}^{\prime}(0)$
$=\left[\frac{1}{2}P_{1}h_{0}(1)-P_{2}h_{0}^{\prime}(1)-\frac{1}{2}P_{1}h_{0}(1)\right]^{\ast}P_{2}^{-1}\left[\frac{1}{2}P_{1}h_{0}(1)+P_{2}h_{0}^{\prime}(1)-\frac{1}{2}P_{1}h_{0}(1)\right]$
$\quad+\left[\frac{1}{2}P_{1}h_{0}(0)-P_{2}h_{0}^{\prime}(0)-\frac{1}{2}P_{1}h_{0}(0)\right]^{\ast}P_{2}^{-1}\left[\frac{1}{2}P_{1}h_{0}(0)+P_{2}h_{0}^{\prime}(0)-\frac{1}{2}P_{1}h_{0}(0)\right]$
$=\begin{bmatrix}-u_{e,0}^{1}+P_{1}y_{e,0}^{1}\\ -u_{e,0}^{2}-P_{1}y_{e,0}^{2}\end{bmatrix}^{\ast}\begin{bmatrix}P_{2}^{-1}&0\\ 0&P_{2}^{-1}\end{bmatrix}\begin{bmatrix}u_{e,0}^{1}-P_{1}y_{e,0}^{1}\\ u_{e,0}^{2}+P_{1}y_{e,0}^{2}\end{bmatrix}$,
Hence,
$\left\|\Big[(2\zeta-1)h_{0}^{\prime}(\zeta)^{\ast}P_{2}h_{0}^{\prime}(\zeta)\Big]_{0}^{1}\right\|\lesssim\|u_{e,0}\|^{2}+\|u_{e,0}\|\|y_{e,0}\|+\|y_{e,0}\|^{2}.$
Thus, the inequality (25) becomes
(26) $\|y_{e,0}\|^{2}$ $\lesssim\|x_{0}\|_{X}^{2}+\frac{r}{|s|}\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}\|x_{0}\|_{X}+\frac{1}{|s|}\|u_{e,0}\|^{2}+\frac{1}{|s|}\|u_{e,0}\|\|y_{e,0}\|$
$\qquad+\frac{1}{|s|}\|y_{e,0}\|^{2}+\frac{1}{|s|}\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}^{2}.$
In order to get an estimate of the transfer function $\mathcal{G}_{e}(s)$, we need to evaluate the norm $\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}$. To this end, we calculate
$\bigg[\bigg(P_{2}$ $h_{0}^{\prime}(\zeta)+\frac{1}{2}P_{1}h_{0}(\zeta)\bigg)^{\ast}P_{2}h_{0}(\zeta)\bigg]^{\prime}$
$=\left(P_{2}h_{0}^{\prime\prime}(\zeta)+\frac{1}{2}P_{1}h_{0}^{\prime}(\zeta)\right)^{\ast}P_{2}h_{0}(\zeta)+\left(P_{2}h_{0}^{\prime}(\zeta)+\frac{1}{2}P_{1}h_{0}(\zeta)\right)^{\ast}P_{2}h_{0}^{\prime}(\zeta).$
Plugging in the first term of equation the equation (21), we get
$=\left(sx_{0}(\zeta)-\frac{1}{2}P_{1}h_{0}^{\prime}(\zeta)\right)^{\ast}P_{2}h_{0}(\zeta)+\|P_{2}h_{0}^{\prime}(\zeta)\|^{2}+\frac{1}{2}h_{0}(\zeta)^{\ast}P_{1}P_{2}h_{0}^{\prime}(\zeta)$
$=\overline{s}x_{0}(\zeta)^{\ast}P_{2}h_{0}(\zeta)-\frac{1}{2}h_{0}^{\prime}(\zeta)^{\ast}P_{1}P_{2}h_{0}(\zeta)-\frac{1}{2}\overline{h_{0}^{\prime}(\zeta)^{\ast}P_{2}P_{1}h_{0}(\zeta)}+\|P_{2}h_{0}^{\prime}(\zeta)\|^{2}.$
This implies
(27) $2\mathrm{Re}$ $\left[\left(P_{2}h_{0}^{\prime}(\zeta)+\frac{1}{2}P_{1}h_{0}(\zeta)\right)^{\ast}P_{2}h_{0}(\zeta)\right]^{\prime}$
$=2\mathrm{Re}\left[\overline{s}x_{0}(\zeta)^{\ast}P_{2}h_{0}(\zeta)\right]+2\|P_{2}h_{0}^{\prime}(\zeta)\|^{2}-h_{0}^{\prime}(\zeta)^{\ast}\left(P_{1}P_{2}+P_{2}P_{1}\right)h_{0}(\zeta).$
Since $P_{2}$ is invertible, we have that the norm $\|P_{2}\cdot\|$ is equivalent to the Euclidean norm $\|\cdot\|$ on $\mathbb{F}^{n}$. Therefore, with (27) we have that
$\int_{0}^{1}\|h_{0}^{\prime}(\zeta)\|^{2}d\zeta$ $\lesssim\int_{0}^{1}\|P_{2}h_{0}^{\prime}(\zeta)\|^{2}d\zeta$
$=\mathrm{Re}\left[\left(P_{2}h_{0}^{\prime}(\zeta)+\frac{1}{2}P_{1}h_{0}(\zeta)\right)^{\ast}P_{2}h_{0}(\zeta)\right]_{0}^{1}$
$\qquad\quad-\int_{0}^{1}\mathrm{Re}[\overline{s}x_{0}(\zeta)^{\ast}P_{2}\mathcal{H}(\zeta)x_{0}(\zeta)]d\zeta$
$\qquad\qquad+\frac{1}{2}\int_{0}^{1}h_{0}(\zeta)^{\ast}\left(P_{1}P_{2}+P_{2}P_{1}\right)h_{0}^{\prime}(\zeta)d\zeta.$
By virtue of $\mathcal{H}\in C^{1}([0,1];\mathbb{F}^{n\times n})$, equation (22), and the Cauchy-Schwarz inequality, we obtain that
$\|h_{0}^{\prime}\|_{\mathrm{L}^{2}}\lesssim\ |s|\|x_{0}\|_{X}^{2}+\|x_{0}\|_{X}\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}+\|u_{e,0}\|\|y_{e,0}\|.$
Therefore, we obtain that for some constant $c>0$
$\|h_{0}^{\prime}\|_{\mathrm{L}^{2}}^{2}-c\|x_{0}\|_{X}\|h_{0}^{\prime}\|_{\mathrm{L}^{2}}-c|s|\|x_{0}\|^{2}_{X}-c\|u_{e,0}\|\|y_{e,0}\|\leq 0.$
This implies
(28) $\|h_{0}^{\prime}\|_{\mathrm{L}^{2}}\leq c\|x_{0}\|_{X}+\sqrt{c^{2}\|x_{0}\|_{X}^{2}+4c\left(|s|\|x_{0}\|_{X}^{2}+\|u_{e,0}\|\|y_{e,0}\|\right)}.$
Since the system is impedance passive, there holds by Corollary 2.3
$\|x_{0}\|_{X}^{2}\lesssim\frac{1}{r}\|u_{e,0}\|\|y_{e,0}\|.$
Thus, for $|s|>1$, the inequality (28) becomes
$\|h_{0}^{\prime}\|_{\mathrm{L}^{2}}$ $\lesssim\frac{1}{\sqrt{r}}\sqrt{\|u_{e,0}\|}\sqrt{\|y_{e,0}\|}+\sqrt{\frac{|s|}{r}\|u_{e,0}\|\|y_{e,0}\|+\|u_{e,0}\|\|y_{e,0}\|}$
$\lesssim\frac{\sqrt{|s|}}{\sqrt{r}}\sqrt{\|u_{e,0}\|}\sqrt{\|y_{e,0}\|}.$
From inequality (26), we recall that
$\|y_{e,0}\|^{2}$ $\lesssim\|x_{0}\|_{X}^{2}+\frac{r}{|s|}\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}\|x_{0}\|_{X}+\frac{1}{|s|}\|u_{e,0}\|^{2}+\frac{1}{|s|}\|u_{e,0}\|\|y_{e,0}\|$
$\qquad\quad+\frac{1}{|s|}\|y_{e,0}\|^{2}+\frac{1}{|s|}\|h_{0}^{\prime}\|_{\mathrm{L^{2}}}^{2}.$
Using again the impedance passivity of the system and the earlier estimate of $\|h_{0}^{\prime}\|_{\mathrm{L}^{2}}$, we deduce that
$\|y_{e,0}\|^{2}$ $\lesssim\frac{1}{r}\|y_{e,0}\|\|u_{e,0}\|+\frac{r}{|s|}\frac{1}{\sqrt{r}}\frac{\sqrt{|s|}}{\sqrt{r}}\|u_{e,0}\|\|y_{e,0}\|+\frac{1}{r}\|u_{e,0}\|^{2}+\frac{1}{r}\|y_{e,0}\|^{2}$
$\qquad\quad+\frac{1}{|s|}\frac{|s|}{r}\|u_{e,0}\|\|y_{e,0}\|$
$\lesssim\left(\frac{2}{r}+\frac{1}{\sqrt{|s|}}\right)\|u_{e,0}\|\|y_{e,0}\|+\frac{1}{r}\|u_{e,0}\|^{2}+\frac{1}{r}\|y_{e,0}\|^{2}.$
Therefore, for some constant $c>0$ (independent of $r$) we have
$\left(1-\frac{c}{r}\right)\|y_{e,0}\|^{2}-\left(\frac{2c}{r}+\frac{c}{\sqrt{|s|}}\right)\|u_{e,0}\|\|y_{e,0}\|-\frac{c}{r}\|u_{e,0}\|^{2}\leq 0.$
For $r>c$, this implies
$\|y_{e,0}\|\leq\ \frac{1}{2\left(1-\frac{c}{r}\right)}\sqrt{c^{2}\left(\frac{2}{r}+\frac{1}{\sqrt{|s|}}\right)^{2}+4\frac{c}{r}\left(1-\frac{c}{r}\right)}\|u_{e,0}\|+\frac{\left(\frac{2}{r}+\frac{1}{\sqrt{|s|}}\right)c}{2(1-\frac{c}{r})}\|u_{e,0}\|.$
Therefore, for $r\gg 1$, we obtain that
(29) $\|y_{e,0}\|\lesssim\ \frac{1}{\sqrt{r}}\|u_{e,0}\|.$
We have shown that for $\mathrm{Re}(s)$ sufficiently large the transfer function satisfies the inequality
(30) $\|\mathcal{G}_{e}(s)\|_{\mathcal{L}(U)}\leq\frac{c}{\sqrt{r}},\quad c>0$
and therefore, it is bounded on a vertical line in the open right half-plane $\mathbb{C}^{+}_{r_{0}}$ for some $r_{0}>0$. By Theorem 5, the system (20) is well-posed. Furthermore, by (30) we obtain that
$\lim_{\mathrm{Re}(s)\to\infty}\mathcal{G}_{e}(s)=0,$
which implies that the system is regular with feedthrough term is zero.
Using the input $u_{e}$ and the output $y_{e}$ defined in Proposition 3.3, see (19) and (20), we may rewrite the system (16). We note that $\begin{bmatrix}u_{e}\\ y_{e}\end{bmatrix}$ is obtained from $\tau(\mathcal{H}x)$ through an invertible linear transformation. Hence $\begin{bmatrix}u_{e}\\ y_{e}\end{bmatrix}$ span the same range as $\tau(\mathcal{H}x)$. Therefore, $u=W_{B,1}\tau(\mathcal{H}x)$, $0=W_{B,2}\tau(\mathcal{H}x)$ and $y=W_{C}\tau(\mathcal{H}x)$ can equivalently be expressed as linear combinations of $u_{e}$ and $y_{e}$, and there exist uniquely determined matrices $K_{1}$, $K_{2}\in\mathbb{F}^{(2n-m)\times 2n}$ and $B_{1}$, $B_{2}$, $C_{1}$, $C_{2}\in\mathbb{F}^{m\times 2n}$ such that the system (16) can equivalently be described by
(31) $\frac{\partial x}{\partial t}(\zeta,t)$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}(\zeta)\right)\mathcal{H}(\zeta)x(\zeta,t),\quad t>0,\quad\zeta\in[0,1]$
(32) $0$ $=K_{1}u_{e}(t)+K_{2}y_{e}(t),$
(33) $u(t)$ $=B_{1}u_{e}(t)+B_{2}y_{e}(t),$
(34) $y(t)$ $=C_{1}u_{e}(t)+C_{2}y_{e}(t).$
The following theorem provides an equivalent condition for the well-posedness of the system (16) in the case $m=2n$. Note that if $m=2n$, then $W_{B,2}=0$ and thus $K_{1}=K_{2}=0$.
Theorem 1.
Let $m=2n$. Then, the boundary control and observation system (16) is well-posed and regular with a feedthrough term given by $C_{1}B_{1}^{-1}$ if and only if the matrix $B_{1}$ in (33) is invertible.
Proof 3.5.
If $B_{1}\in\mathbb{F}^{2n\times 2n}$ is invertible, then the equation (33) may be reformulated as
$u_{e}(t)=B_{1}^{-1}u(t)-B_{1}^{-1}B_{2}y_{e}(t).$
Thus, the system (16) is the system (20) closed via the output feedback defined by $B_{1}^{-1}B_{2}$. From Proposition 3.3, we have $\lim_{\mathrm{Re}\,s\to\infty}\mathcal{G}_{e}(s)=0$, then, for $\mathrm{Re}(s)$ large enough we get that $I+B_{1}^{-1}B_{2}\mathcal{G}_{e}(s)$ is invertible and its inverse exists and is bounded on a right half-plane. Therefore, the matrix $B_{1}^{-1}B_{2}$ defines an admissible feedback operator and thus by Theorem 7, the closed loop system is a well-posed boundary control and observation system with a transfer function is given by
$\mathcal{G}(s)=\left(C_{1}+C_{2}\mathcal{G}_{e}(s)\right)\left(I+B_{1}^{-1}B_{2}\mathcal{G}_{e}(s)\right)^{-1},\quad s\in C^{+}_{\alpha}$
for some $\alpha>0$. Furthermore, we see that it converges to $C_{1}B_{1}^{-1}$, as $\mathrm{Re}(s)\to\infty$.
Conversely, we assume that the system (16) with $m=2n$ is well-posed and we show that $B_{1}$ is invertible. By assumption, the matrix $\begin{bmatrix}W_{B,1}\\ W_{C}\end{bmatrix}$ is invertible. Hence, we conclude that $\begin{bmatrix}B_{1}&B_{2}\\ C_{1}&C_{2}\end{bmatrix}$ is invertible as well. As the system (20) is well-posed with a feedthrough 0, we apply Lemma 2.6 and obtain that $B_{1}$ is invertible.
If $0<m\leq 2n$, the sufficient condition of Theorem 1 still holds, which will be shown in the following theorem. Further, we provide a necessary condition.
Corollary 2.
The boundary control and observation system (16) is well-posed if the matrix
$\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}=\begin{bmatrix}W_{B,1}\\ W_{B,2}\end{bmatrix}P_{2}^{-1}\begin{bmatrix}0&0\\ I&0\\ 0&0\\ 0&I\end{bmatrix}$
in (33) is invertible.
Proof 3.6.
If the matrix $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ in (33) is invertible, the well-posedness of the system (16) follows from the fact that we may introduce the additional input $v(t)=W_{B,2}\mathcal{H}\tau(x)$. The system with the extended input $\begin{bmatrix}u\\ v\end{bmatrix}$ is well-posed by Theorem 1 as $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ is invertible. Hence we may choose any input that is (locally) square integrable. In particular, we can choose $v\equiv 0$ and thus the boundary control and observation system (16) is well-posed if the matrix $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ in (33) is invertible.
4 Exact controllability and exact observability
After characterizing the well-posedness of the class of boundary control and observation systems (16) in the previous section, we now turn our attention to the property of exact controllability. Throughout this section we assume that $P_{0}^{\ast}(\zeta)=-P_{0}(\zeta)$ for a.e. $\zeta\in(0,1)$ and $m=2n$. Under this assumption, the system takes the form
(35) $\frac{\partial x}{\partial t}(\zeta,t)$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}(\zeta)\right)\mathcal{H}(\zeta)x(\zeta,t),\quad t>0,\quad\zeta\in[0,1]$
$u(t)$ $=W_{B,1}\tau(\mathcal{H}x)(t),\quad t>0$
$y(t)$ $=W_{C}\tau(\mathcal{H}x)(t),\quad t>0$
$x(\zeta,0)$ $=x_{0}(\zeta),\quad\zeta\in[0,1].$
From this point onward, we define
$\mathcal{A}x$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}\right)\mathcal{H}x,$
$\mathcal{B}x$ $=W_{B,1}\tau(\mathcal{H}x),$
$\mathcal{C}x$ $=W_{C}\tau(\mathcal{H}x),$
$D(\mathcal{A})$ $=\left\{x\in\mathrm{L}^{2}((0,1);\mathbb{F}^{n})\,|\,\mathcal{H}x\in\mathrm{H}^{2}((0,1);\mathbb{F}^{n})\right\},$
$D(\mathcal{B})$ $=D(\mathcal{A}),$
and we assume that the system (35) is a well-posed boundary control and observation system and we aim to characterize the exact controllability and exact observability in finite time.
Definition 1.
[TucsnakWeiss:14, Definition 6.2.1 and Definition 6.2.12] The boundary control and observation system (3) is called exactly controllable (in finite time), if there exists a time $T>0$ such that for every $x_{1}\in X$ there exists a control function $u\in\mathrm{L^{2}}((0,T);\mathbb{F}^{m})$ such that the corresponding mild solution $x$ with $x_{0}=0$ satisfies
$x(0)=0,\quad x(T)=x_{1}.$
It is called exactly observable (in finite time) if there exists a time $T>0$ and a constant $m>0$ such that for every initial condition $x_{0}\in X$ the corresponding mild solution $(x,y)$ (with $u=0$) satisfies
$\int_{0}^{T}\|y(t)\|^{2}dt\geq m\|x_{0}\|^{2}.$
We start by providing the definition of optimizability.
Definition 2.
[Weiss_Rebarber:97, Definition 1.1] A boundary control and observation system is called optimizable if for every initial condition $x_{0}\in X$, there exists an input function $u\in\mathrm{L^{2}}((0,\infty);\mathbb{F}^{m})$ such that the mild solution $x$ satisfies
$\int_{0}^{\infty}\|x(t)\|^{2}dt<\infty.$
Note that exact controllability implies optimizability and in [Weiss_Rebarber:97], a sufficient condition for exact controllability is presented in terms of optimizability.
Proposition 4.1.
[Weiss_Rebarber:97, Corollary 2.2] If the boundary control and observation system (3) is optimizable and $A_{0}\coloneqq-\mathcal{A}_{\mid\ker\mathcal{B}}$ generates a bounded $C_{0}$-semigroup, then the system (3) is exactly controllable.
For boundary control and observation system (35) which conserves energy over time, are naturally exactly controllable.
Lemma 4.2.
Assume that the boundary control and observation system (35) is impedance energy preserving, then it is exactly controllable.
Proof 4.3.
We consider the operator
$A_{0}x$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}\right)\mathcal{H}x,$
$D(A_{0})$ $=\left\{x\in X\,|\ \mathcal{H}x\in\mathrm{H^{2}}((0,1);\mathbb{F}^{n}),\ W_{B,1}\tau(\mathcal{H}x)=0\right\}.$
As the boundary control system is impedance energy-preserving, the operator $A_{0}\coloneqq\mathcal{A}_{\mid\ker\mathcal{B}}$ generates a strongly continuous unitary group [GorZwaMas:2005]. Thus, $A_{0}$ generates a strongly continuous bounded semigroup. Therefore, by Proposition 4.1 exact controllability is equivalent to optimizability. Accordingly, we show that the system can be exponentially stabilized by a negative output feedback, that is u(t)=ky(t),k>0u(t)=-ky(t),\,k>0. Indeed, for $x_{0}\in D(\mathcal{A})\cap\ker\mathcal{B}+k\mathcal{C}$
(36) $\frac{d}{dt}\left\|T(t)x_{0}\right\|$ $=2\mathrm{Re}\left\langle\mathcal{A}x(t),x(t)\right\rangle_{X}$
$=2\mathrm{Re}\left\langle u(t),y(t)\right\rangle_{U}$
$=2\mathrm{Re}\left\langle-ky(t),y(t)\right\rangle_{U}$
$=-2k\left\|y(t)\right\|^{2}.$
Let $t_{f}$ be such that $\left\|T(t_{f})x_{0}\right\|\leq\frac{1}{2}\left\|x_{0}\right\|$. By integrating the equation (36) over the interval $0,t_{f}]$, we obtain
$\left\|T(t_{f})x_{0}\right\|-\left\|x_{0}\right\|=-2k\int_{0}^{t_{f}}\left\|y(t)\right\|^{2}dt$
Therefore, we obtain
$\int_{0}^{t_{f}}\left\|y(t)\right\|^{2}dt\geq m_{t_{f}}\left\|x_{0}\right\|,\quad m_{t_{f}}>0.$
This shows that the system closed by the negative output feedback is exactly observable. Thus, the semigroup generated by $\mathrm{A}_{k}:=\mathcal{A}_{\mid\ker\mathcal{B}+k\mathcal{C}}$ is exponentially stable. This leads to a mild solution in $\mathrm{L^{2}}((0,\infty);X)$. Hence, the system is optimizable and thus it is exactly controllable.
In the light of duality concept between controllability and observability [tucsnak2009observation], we next define the dual system of (35) to derive an analogous result for exact observability.
Lemma 4.4.
[GorZwaMas:2005, Lemma A.1] A matrix $W\in\mathbb{F}^{2n\times 4n}$ has full row rank if and only if $W$ has the decomposition
$W=S\begin{bmatrix}I+V&I-V\end{bmatrix}$,
where $S,V\in\mathbb{F}^{2n\times 2n}$ such that $S$ is invertible. Furthermore, $W\Sigma W^{\ast}\geq 0$ (resp. $W\Sigma W^{\ast}=0$) if and only if $VV^{\ast}\leq I$ (resp. $VV^{\ast}=I$), where $\Sigma=\begin{bmatrix}0&I\\ I&0\end{bmatrix}$.
Lemma 4.5.
[Villegas:07, Theorem 2.24] If the boundary control and observation system (35) is impedance passive, then the operators $(\mathcal{A}^{\ast},\mathcal{B}^{\ast},\mathcal{C}^{\ast})$ describing the dual system of (35) are given by
$\mathcal{A}^{\ast}x$ $=\left(-P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}-P_{0}\right)\mathcal{H}x,$
$D(\mathcal{A}^{\ast})$ $=\left\{x\in X\,|\,\mathcal{H}x\in\mathrm{H}^{2}((0,1);\mathbb{F}^{n})\right\},$
$\mathcal{B}^{\ast}x(t)$ $=\begin{bmatrix}I-V^{\ast}&-I-V^{\ast}\end{bmatrix}\begin{bmatrix}R^{-1}&-R^{-1}\\ I&I\end{bmatrix}\tau(\mathcal{H}x),$
$D(\mathcal{B}^{\ast})$ $=D(\mathcal{A}^{\ast}),$
$\mathcal{C}^{\ast}x(t)$ $=W_{C}\tau(\mathcal{H}x),$
where
$\tilde{W}_{B}=W_{B,1}\begin{bmatrix}R&I\\ -R&I\end{bmatrix}=S\begin{bmatrix}I+V&I-V\end{bmatrix}$,
and
$R=\left[\begin{matrix}0&-P_{2}^{-1}\\ P_{2}^{-1}&P_{2}^{-1}P_{1}P_{2}^{-1}\end{matrix}\right],$
with $S,V\in\mathbb{F}^{2n\times 2n}$, $VV^{\ast}\leq I$ and $S$ invertible.
Lemma 4.6.
If the boundary control and observation system (35) is impedance energy preserving, then it is exactly observable.
Proof 4.7.
It is worth noting that the dual system also satisfies the conditions of Assumption 3.1. Furthermore, the system (35) is impedance energy preserving implies that the dual system is also impedance energy preserving. Indeed, by Remark 3.2, the system is impedance energy preserving is equivalent to $\tilde{W}_{B}\Sigma\tilde{W}_{B}^{\ast}=0$. Therefore, by Lemma 4.4, there exists $S,~V\in\mathbb{F}^{2n\times 2n}$ with $S$ invertible and $VV^{\ast}=I$ such that
$\tilde{W}_{B}=S\begin{bmatrix}I+V&I-V\end{bmatrix}$.
Furthermore, we obtain that
$\begin{bmatrix}I-V^{\ast}&-I-V^{\ast}\end{bmatrix}\Sigma\begin{bmatrix}I-V\\ -I-V\end{bmatrix}=0.$
By Lemma 4.5, we observe that the dual system of (35) is indeed an impedance energy preserving system. Therefore, using Lemma 4.2, we conclude that the dual system is exactly controllable. Since controllability and observability are dual notions [tucsnak2009observation], the system (35) is exactly observable.
A key observation in this framework is the use of the following proposition, which states that given a well-posed boundary control and observation system, the exact controllability and exact observability remains unchanged under admissible feedback.
Proposition 4.8.
[Wei:94, Remark 6.9] Let $F$ is an admissible feedback operator for the well-posed boundary control and observation system (35). Then the closed-loop boundary control and observation system generated by the operators $(\mathcal{A},(\mathcal{B}-F\mathcal{C}),\mathcal{C})$ is exactly controllable (resp. exactly observable) if and only if the open-loop system generated by the operators $(\mathcal{A},\mathcal{B},\mathcal{C})$ is exactly controllable (resp. exactly observable).
The following theorem provides a sufficient condition for exact controllability and exact observability of the system (35).
Theorem 3.
The well-posed boundary control and observation system (35) is exactly controllable and exactly observable.
Proof 4.9.
By Proposition 3.3, the boundary control and observation system
(37) $\frac{\partial x}{\partial t}(\zeta,t)$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}(\zeta)\right)\mathcal{H}(\zeta)x(\zeta,t),\quad\zeta\in[0,1]$
$u_{e}(t)$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}P_{2}\frac{\partial}{\partial\zeta}(\mathcal{H}x)(1,t)+\tfrac{1}{2}P_{1}(\mathcal{H}x)(1,t)\\ P_{2}\frac{\partial}{\partial\zeta}(\mathcal{H}x)(0,t)+\tfrac{1}{2}P_{1}(\mathcal{H}x)(0,t)\end{bmatrix}$,
$y_{e}(t)$ $=\frac{1}{\sqrt{2}}\begin{bmatrix}(\mathcal{H}x)(1,t)\\ -(\mathcal{H}x)(0,t)\end{bmatrix}$,
is impedance energy preserving. Then, it follows from Lemma 4.2 (resp. Lemma 4.6) that it is exactly controllable (resp. exact observable). Since the system (35) is well-posed, then $B^{-1}$ is invertible. Using
$u(t)$ $=B_{1}u_{e}(t)+B_{2}y_{e}(t),$
$y(t)$ $=C_{1}u_{e}(t)+C_{2}y_{e}(t),$
we see that the system (35) is the closed-loop system of (37) via the feedback defined by $B_{1}^{-1}B_{2}$. Since the feedthrough term of the system (37) is zero, then $B_{1}^{-1}B_{2}$ is an admissible feedback operator. Thus, from Lemma 4.8, we deduce that the closed-loop system, that is the system
$\frac{\partial x}{\partial t}(\zeta,t)$ $=\left(P_{2}\frac{\partial^{2}}{\partial\zeta^{2}}+P_{1}\frac{\partial}{\partial\zeta}+P_{0}(\zeta)\right)\mathcal{H}(\zeta)x(\zeta,t),\quad t>0,\quad\zeta\in[0,1]$
$u(t)$ $=W_{B,1}\tau(\mathcal{H}x),\quad t>0$
$y(t)$ $=W_{C}\tau(\mathcal{H}x),\quad t>0.$
is exactly controllable and exactly observable.
5 Examples
As an illustrative example, we apply this theory to the Euler-Bernoulli models.
5.1 Euler-Bernoulli beam with viscous air damping
We study an Euler-Bernoulli beam with viscous air damping. In this model, damping is assumed proportional to the transverse velocity, $\frac{\partial\omega}{\partial t}$ [BanksInman_91]. In this case the vibrations of the beam are described by
(38) $\rho(\zeta)\frac{\partial^{2}\omega}{\partial t^{2}}(\zeta,t)+\frac{\partial^{2}}{\partial\zeta^{2}}\left(EI(\zeta)\frac{\partial^{2}\omega}{\partial\zeta^{2}}(\zeta,t)\right)+\gamma(\zeta)\frac{\partial\omega}{\partial t}(\zeta,t)=0,\quad t>0,\quad\zeta\in(0,1).$
Here, $\omega(\zeta,t)$ is the vertical displacement, $\rho(\zeta)$ denotes the mass density per unit length, $EI(\zeta)>0$ is the flexural rigidity (or bending stiffness) of the beam while $E(\zeta)>0$ is the elasticity modulus and $I(\zeta)>0$ is the moment of inertia of the cross section. The non-negative coefficient $\gamma(\zeta)$ represents the viscous external damping. Associated with (38) we choose the following boundary conditions
$\frac{\partial^{2}\omega}{\partial t\partial\zeta}(0,t)$ $=0,\quad\frac{\partial^{2}\omega}{\partial t\partial\zeta}(1,t)=0,$
$\left[\frac{\partial}{\partial\zeta}\left(EI(\zeta)\frac{\partial^{2}\omega}{\partial\zeta^{2}}(\zeta,t)\right)\right]_{\zeta=0}$ $=0,\quad\left[\frac{\partial}{\partial\zeta}\left(EI(\zeta)\frac{\partial^{2}\omega}{\partial\zeta^{2}}(\zeta,t)\right)\right]_{\zeta=1}=u(t),$
and measurement of the output
$y(t)=\frac{\partial\omega}{\partial t}(1,t).$
By defining $x=\begin{bmatrix}\rho\frac{\partial\omega}{\partial t}\\ \frac{\partial^{2}\omega}{\partial\zeta^{2}}\end{bmatrix}$, and choosing
$P_{2}=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix},\quad P_{1}=0,\quad P_{0}=\begin{bmatrix}-\gamma&0\\ 0&0\end{bmatrix},\text{ and}\quad\mathcal{H}=\begin{bmatrix}\frac{1}{\rho}&0\\ 0&EI\end{bmatrix}$,
the damped Euler-Bernoulli beam equations (38) may be written as the class of system (16) with the following boundary control and observation
$\begin{bmatrix}0\\ I\end{bmatrix}u(t)$ $=\begin{bmatrix}0&0&0&0&0&0&1&0\\ 0&0&1&0&0&0&0&0\\ 0&0&0&0&0&0&0&1\\ 0&0&0&1&0&0&0&0\\ \end{bmatrix}\begin{bmatrix}(\mathcal{H}x)(1,t)\\ (\mathcal{H}x)^{\prime}(1,t)\\ (\mathcal{H}x)(0,t)\\ (\mathcal{H}x)^{\prime}(0,t)\end{bmatrix}$,
$y(t)$ $=\begin{bmatrix}1&0&0&0&0&0&0&0\end{bmatrix}\begin{bmatrix}(\mathcal{H}x)(1,t)\\ (\mathcal{H}x)^{\prime}(1,t)\\ (\mathcal{H}x)(0,t)\\ (\mathcal{H}x)^{\prime}(0,t)\end{bmatrix}$.
We have
$\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}=\begin{bmatrix}0&0&0&1\\ 0&1&0&0\\ 0&0&-1&0\\ -1&0&0&0\end{bmatrix}$.
We clearly have that $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ is invertible. Thus, by Corollary 2, with this choice of input and output, the system is well-posed.
5.2 Beam with distributed elastic support
We consider an Euler-Bernoulli beam where its end are supported transversely by distributed elastic springs where $k_{t}$ denotes the translational stiffness and $k_{r}$ denotes the rotational stiffness of the spring, see [mathews1958vibrations, ElasticBeam:19]. The equation of motion for this system is given by
(39) $\rho(\zeta)\frac{\partial^{2}\omega}{\partial t^{2}}(\zeta,t)+\frac{\partial^{2}}{\partial\zeta^{2}}\left(EI(\zeta)\frac{\partial^{2}\omega}{\partial\zeta^{2}}(\zeta,t)\right)+k\omega(\zeta,t)=0,\quad t>0,\quad\zeta\in(0,1).$
The Euler-Bernoulli beam equation (39) may be written in the formulation of the class of system (16)
$\frac{\partial}{\partial t}\begin{bmatrix}\rho\frac{\partial\omega}{\partial t}\\ \frac{\partial^{2}\omega}{\partial\zeta^{2}}\\ \omega\end{bmatrix}=\begin{bmatrix}0&-1&0\\ 1&0&0\\ 0&0&i\end{bmatrix}\frac{\partial^{2}}{\partial\zeta^{2}}\begin{bmatrix}\frac{1}{\rho}&0&0\\ 0&EI&0\\ 0&0&k\end{bmatrix}\begin{bmatrix}\rho\frac{\partial\omega}{\partial t}\\ \frac{\partial^{2}\omega}{\partial\zeta^{2}}\\ \omega\end{bmatrix}+\begin{bmatrix}0&0&-1\\ 0&0&0\\ 1&\frac{-ik}{EI}&0\end{bmatrix}\begin{bmatrix}\frac{1}{\rho}&0&0\\ 0&EI&0\\ 0&0&k\end{bmatrix}\begin{bmatrix}\rho\frac{\partial\omega}{\partial t}\\ \frac{\partial^{2}\omega}{\partial\zeta^{2}}\\ \omega\end{bmatrix}$.
We choose the following boundary conditions
$EI(0)\frac{\partial^{2}\omega}{\partial\zeta^{2}}(0,t)+k_{r}\frac{\partial\omega}{\partial\zeta}(0,t)$ $=0,$
$\frac{\partial^{2}\omega}{\partial t\partial\zeta}(0,t)$ $=0,$
$\left[\frac{\partial}{\partial\zeta}\left(EI(\zeta)\frac{\partial^{2}\omega}{\partial\zeta^{2}}(\zeta,t)\right)\right]_{\zeta=0}+k_{t}w(0,t)$ $=0,$
$\frac{\partial^{2}\omega}{\partial t\partial\zeta}(1,t)$ $=0,$
$EI(1)\frac{\partial^{2}\omega}{\partial\zeta^{2}}(1,t)+k_{r}\frac{\partial\omega}{\partial\zeta}(1,t)$ $=0,$
$\left[\frac{\partial}{\partial\zeta}\left(EI(\zeta)\frac{\partial^{2}\omega}{\partial\zeta^{2}}(\zeta,t)\right)\right]_{\zeta=1}+k_{t}\omega(1,t)$ $=u(t).$
We introduce the output
$y(t)=\frac{\partial\omega}{\partial t}(1,t).$
We have that
$\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}=\begin{bmatrix}0&0&0&0&0&-ik_{r}\\ 0&0&0&0&1&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ 0&0&-ik_{r}&0&0&0\\ 1&0&0&0&0&0\end{bmatrix}$.
Since $\begin{bmatrix}K_{1}\\ B_{1}\end{bmatrix}$ is invertible, by Corollary 2, the system is well-posed.
6 Conclusion and perspectives
In this work, we have investigated the well-posedness of a class of boundary control and observation system in the form (16) which in particular include the Euler-Bernoulli beam models. To this end, we have provided a necessary and sufficient condition for the well-posedness. As a consequence, we have showed that, under full control and observation, well-posedness implies exact controllability. A perspective could be to extend the proposed approach to a more general class that include the Rayleigh beam equation.